# Thought Anchors: Which LLM Reasoning Steps Matter?

### v3

@@ D.2 Attention aggregation
The R1-Distill-Qwen-14B and R1-Distill-Llama-8B receiver heads displayed consistent patterns related to sentence types, such that plan generation, uncertainty management, and self checking sentences received heightened attention; although visually, the differences to fact retrieval and active computation may be less prominent, paired t-tests (paired with respect to a given response) showed that plan generation, uncertainty management, and self checking always significantly surpassed fact retrieval and active computation (six paired t-tests: $p$s $\leq.01$).
@@
Figure 11: The plots here show the vertical-attention score patterns associated with the R1-Distill-Llama-8B data. A. This histogram shows the kurtosis values across all attention heads, averaged across all reasoning traces; parallels Figure 4 based on the R1-Qwen-14B data. B. This scatterplot shows the kurtosis of each head’s vertical-attention score, organized by layer. Figure 17 is the R1-Distill-Qwen-14B version of this figure, which showed an upward trend into later layers that is not evident here.
@@
Appendix E Embeddings model
We used all-MiniLM-L6-v2 with a maximum sequence length of $256$ tokens and a hidden dimension of $384$ as our sentence embeddings model from the sentence-transformers [Reimers and Gurevych, 2019] library. We picked a cosine similarity threshold of $0.8$, which is the median similarity value between all sentence removed (i.e., original sentence) and sentence resampled pairs in our dataset.
Appendix F Additional resampling results
Appendix G Additional resampling details
G.1 Counterfactual versus resampling importance
G.2 Sentence-to-sentence importance
Output: Importance score importance(i -> j)
1. Get rollouts R_keep where sentence i was kept (resampling from i+1)
3. For each rollout r in R_keep:
   a. Extract all sentences S_r from rollout r
   b. Find best matching sentence to target sentence j:
      - Compute sentence embeddings
      - Calculate cosine similarity between each s in S_r and target j
      - Select sentence with highest similarity if similarity >= t
   c. Add to matches_keep if valid match found
4. For each rollout r in R_remove:
   a. Extract all sentences S_r from rollout r
   c. Add to matches_remove if valid match found
5. Calculate match rates:
   match_rate_keep = |matches_keep| / |R_keep|
   match_rate_remove = |matches_remove| / |R_remove|
6. Return importance(i -> j) = match_rate_keep - match_rate_remove
Appendix H Additional receiver head information
Figure 18: The attention weight matrices for the receiver head with the highest kurtosis score are shown here for all twenty responses. The coloring was defined such that the darkest navy corresponds to values surpassing 99.5th percentile value of each matrix. White is zero.
Appendix I Reasoning versus base model differences in receiver heads
Appendix J Effects of ablating receiver heads
J.1 Methodology
  1. 1.
  2. 2.
J.2 Results and Discussion
Heads Ablated Receiver heads Random heads
256 48.8% [39.3%, 58.3%] 52.7% [43.0%, 62.5%]
512 27.7% [17.2%, 38.2%] 37.3% [27.5%, 47.1%]
Appendix K Receiver head correlations with sentence-sentence resampling importance
A sentence’s receiver-head score predicts its impact on downstream sentences per the resampling method. Specifically, our resampling approach (Section˜3) produces a matrix describing each sentence’s counterfactual importance on each future sentence. Based on said matrix, we can compute a sentence-level measure analogous to the receiver head score by taking the average value below the matrix diagonal; as with the receiver head score, we omit the four entries immediately below the diagonal to avoid simply capturing proximal effects and focus on farther downstream impacts; we take the absolute value of the resampling matrix to produce an estimate of how a given sentence may both upregulate or downregulate future sentences.
We correlated sentence (top-64-head) receiver scores with their resampling scores calculated here - correlated separately for each response - yielding a positive association (mean Spearman $\rho=.22$; 95% CI: [.17, .30]). Although the correlation is nominally weak/moderate, this is a consistent trend, and 18 of the 20 responses tested displayed a positive association ($\rho>0$). A sentence’s receiver-score additionally correlates with its downstream effect as computed by the attention suppression measure introduced in Section˜5 (mean >4 below the diagonal) (mean Spearman $\rho=.19$; 95% CI: [.10, .28]). Although attention weights alone are not a causal measure, they are computationally cheap to examine, requiring no post hoc forward pass, unlike the other two methods. Grounded in these other techniques, the receiver-head scores may hence prove useful by virtue of their limited cost.
Appendix L Case study transcript
Below is a transcript of the model’s response used for the case study (Section˜6). The first 74 sentences following <think> are shown. After this point, the model converged on 100% resampling accuracy for all subsequent sentences.
@@ Appendix I Reasoning versus base model differences in receiver heads
Appendix M Sentence position effects on receiver-head scores

### v4

@@ B.2 Attention aggregation
The R1-Distill-Qwen-14B and R1-Distill-Llama-8B receiver heads displayed consistent patterns related to sentence types, such that plan generation, uncertainty management, and self checking sentences received heightened attention; although visually, the differences to fact retrieval and active computation may be less prominent, paired t-tests (paired with respect to a given response) showed that plan generation and uncertainty management always significantly surpassed fact retrieval and active computation (four paired t-tests: $p$s $\leq.01$).
@@
Figure 11: The plots here show the vertical-attention score patterns associated with the R1-Distill-Llama-8B data. A. This histogram shows the kurtosis values across all attention heads, median across all reasoning traces; parallels Figure 4 based on the R1-Qwen-14B data. B. This scatterplot shows the kurtosis of each head’s vertical-attention score, organized by layer. Figure 19 is the R1-Distill-Qwen-14B version of this figure, which showed an upward trend into later layers that is not evident here.
@@
Appendix C Case study details
C.1 Case study transcript
Below is a transcript of the model’s response used for the case study (Section˜5.2). The first 74 sentences following <think> are shown. After this point, the model converged on 100% resampling accuracy for all subsequent sentences.
@@ C.2 Case study: Sentence 13 alternatives
C.3 Receiver head and sentence-sentence case study findings
The presented techniques cover different aspects of attribution within a reasoning trace. Building on the case-study conclusions from our resampling approach (section˜2.4), we study the model’s CoT here by focusing on receiver heads and sentence-sentence links (Figure 13) (see above, Appendix˜C, for the full transcript).
C.3.1 Receiver heads
C.3.2 Attention suppression
C.4 Sentence position effects on receiver-head scores
@@
Appendix D Sentence taxonomy
Building on top of the framework presented by (Venhoff et al., 2025), we developed a taxonomy consisting of eight distinct sentence categories that capture reasoning functions in mathematical problem-solving. Each category represents a specific cognitive operation. The functions and examples for each category are given in Table 1. Notably, the uncertainty management category includes backtracking sentences.
Category Function Examples
Appendix E Prompt information
Function Tags:
Dependencies:
- If there’s no clear dependency use an empty list: []
Important Notes:
- Make sure to include all dependencies for each sentence.
- Include both long-range and short-range dependencies.
- Do NOT forget about long-range dependencies.
- Try to be as comprehensive as possible.
- Make sure there is a path from earlier sentences to the final answer.
Output Format:
Return a dictionary with one entry per sentence, where each entry has:
- the sentence index (as the key, converted to a string),
- a dictionary with:
- "function_tags": list of tag strings
- "depends_on": list of sentence indices, converted to strings
Here is the expected format:
"1": {
"function_tags": ["problem_setup"],
"depends_on": [""]
"4": {
"function_tags": ["plan_generation"],
"depends_on": ["3"]
"5": {
"function_tags": ["fact_retrieval"],
"depends_on": []
"9": {
"function_tags": ["active_computation"],
"depends_on": ["4", "5"]
"24": {
"function_tags": ["uncertainty_management"],
"depends_on": ["23"]
"32": {
"function_tags": ["final_answer_emission"],
"depends_on": ["9, "30", "32"]
Here is the math problem:
<PROBLEM>
Here is the full chain-of-thought, broken into sentences:
<SENTENCES>
Now label each sentence with function tags and dependencies.
Appendix F Sentence category probing
Appendix G Embeddings model
We used all-MiniLM-L6-v2 with a maximum sequence length of $256$ tokens and a hidden dimension of $384$ as our sentence embeddings model from the sentence-transformers (Reimers & Gurevych, 2019) library. We picked a cosine similarity threshold of $0.8$, which is the median similarity value between all sentence removed (i.e., original sentence) and sentence resampled pairs in our dataset.
Appendix H Additional resampling results
Appendix I Counterfactual versus resampling importance
Appendix J Additional receiver head information
Figure 20: The attention weight matrices for the receiver head with the highest kurtosis score are shown here for twenty of the forty responses (selected arbitrarily based on the first twenty processed). The coloring was defined such that the darkest navy corresponds to values surpassing 99.5th percentile value of each matrix. White is zero.
Appendix K Reasoning versus base model differences in receiver heads
Appendix L Effects of ablating receiver heads
L.1 Methodology
  1. 1.
  2. 2.
L.2 Results and Discussion
Heads Ablated Receiver heads Random heads
256 48.8% [39.3%, 58.3%] 52.7% [43.0%, 62.5%]
512 27.7% [17.2%, 38.2%] 37.3% [27.5%, 47.1%]
Appendix M KL causal graph pseudocode
This pseudocode outlines the procedure for computing a sentence-to-sentence causal graph for a given chain-of-thought (CoT). The algorithm works by systematically masking each source sentence and measuring the resulting change in the model’s predictions (logits) for all subsequent target sentences. The sentence–sentence impact is quantified as the average log-KL divergence across a target sentence’s tokens, which is then normalized against the average impact from all prior sentences. This last normalization step effectively accounts for differences in target sentences’ average entropy, which may vary widely and can hamper studying differences between target sentences.
Masking can be performed either by suppressing attention toward the source sentence or omitting the sentence entirely; the former preserves positional embedding information, while the latter may be computationally cheaper and easier to implement (e.g., with serverless providers). If masking is done by omitting sentence $i$ from the CoT, rather than suppressing attention toward sentence $i$, this will impact sentence $j$’s token positions across the CoT and masked CoT, which should be accounted for.
Algorithm 1 GetCausalMatrix(CoT, Model)
1:Initialize $\textsc{Causal\_Matrix}\in\mathbb{R}^{M\times M}\leftarrow 0$ $\triangleright$ $M$ = number of sentences in CoT
[…]
3:for $i=0$ to M1M-1 do $\triangleright$ source sentence
4:  $\text{CoT}_{\text{masked}}\leftarrow\textsc{Mask\_Source}(\text{CoT},i)$
5:  $\textsc{logits\_masked}\leftarrow\textsc{Forward\_Pass}(\text{CoT}_{\text{masked}},\text{Model})$
6:  for $j=i+1$ to M1M-1 do $\triangleright$ target sentence
7:   $\textsc{tokens\_j}\leftarrow\textsc{Sentence\_Tokens}(\text{CoT}_{\text{masked}},j)$
8:   $\textsc{total\_KL}\leftarrow 0$
9:   for each $k\in\textsc{tokens\_j}$ do
10:     $\textsc{KL}\leftarrow\textsc{KLDivergence}(\textsc{logits\_base}[k],\textsc{logits\_masked}[k])$
11:     $\textsc{total\_KL}\leftarrow\textsc{total\_KL}+\log(\textsc{KL})$
12:   end for
13:   $\textsc{Causal\_Matrix}[i,j]\leftarrow\textsc{total\_KL}/|\textsc{tokens\_j}|$
14:  end for
15:end for
[…]
17:for $j=0$ to M1M-1 do
18:  $\mu\leftarrow\textsc{mean}(\textsc{Causal\_Matrix}[0\!:\!j,\,j])$
19:  $\textsc{Causal\_Matrix}[0\!:\!j,\,j]\leftarrow\textsc{Causal\_Matrix}[0\!:\!j,\,j]-\mu$
20:end for
21:return Causal_Matrix
[…]
5:for each vocabulary index $v$ do
6:  $\textsc{kl}\leftarrow\textsc{kl}+p[v]\cdot\big(\log p[v]-\log q[v]\big)$
7:end for
8:return kl
Appendix N Sentence-to-sentence counterfactual importance
We extend our counterfactual resampling framework (section˜3.1) to quantify each sentence’s influence on each future sentence. Further below, we describe how this measure’s values for sentence-sentence links correlate with the values generated via our section˜5 method, masking sentences and measuring the impact on later sentences’ logits.
We estimate the counterfactual importance of sentence $S_{i}$ on a future sentence $S^{\text{Fut.}}$ formally with:
In this context $S^{\text{Fut.}}$ semantically occurs if, when we extract the sentences and identify the best candidate match for $S^{\text{Fut.}}$ using cosine similarity between sentence embeddings, it has greater than $0.8$ cosine similarity (i.e., the median value in our dataset) to that sentence. Pseudocode for estimating sentence-to-sentence importance and empirical values of this metric can be found in Appendix˜I.
Output: Importance score importance(i -> j)
1. Get rollouts R_keep where sentence i was kept (resampling from i+1)
3. For each rollout r in R_keep:
   a. Extract all sentences S_r from rollout r
   b. Find best matching sentence to target sentence j:
      - Compute sentence embeddings
      - Calculate cosine similarity between each s in S_r and target j
      - Select sentence with highest similarity if similarity >= t
   c. Add to matches_keep if valid match found
4. For each rollout r in R_remove:
   a. Extract all sentences S_r from rollout r
   c. Add to matches_remove if valid match found
5. Calculate match rates:
   match_rate_keep = |matches_keep| / |R_keep|
   match_rate_remove = |matches_remove| / |R_remove|
6. Return importance(i -> j) = match_rate_keep - match_rate_remove
N.1 Correlations with the resampling-based importance matrix
The attention-suppression matrix values correlate with those of the resampling-method matrix. Specifically, the two matrices were positively correlated for 90% of reasoning traces (mean $r=.20$, 95% CI: [.12, .27]). This association is stronger when considering only cases fewer than five sentences apart in the reasoning trace, which may better track direct rather than indirect effects represented by the resampling method (mean $r=.34$ [.27, .40]). The magnitudes of these correlations are substantial, given that the two measures capture partially different aspects of causality and the resampling measure itself contains stochastic noise. Hence, these results give weight to the validity of the resampling approach, whose precision we leverage for the forthcoming case study.